# Enhanced entanglement scaling and area-law charge fluctuations in a non-Fermi liquid of composite fermions

### v1

@@ Abstract
The composite fermion Fermi liquid (CFL) state at $\nu{=}1/2$ filling of a Landau level is a paradigmatic example of a non-Fermi liquid borne out purely by Coulomb interactions. But in what ways is this exotic state of matter precisely different from a Fermi liquid? The entanglement entropy of the CFL state was indeed found to exhibit a significant enhancement compared to free electrons [Shao et al., Phys. Rev. Lett. 114, 206402 (2015)], which was subsequently ruled out as a finite-size effect by the study of a lattice CFL analogue [Mishmash and Motrunich, Phys. Rev. B 94, 081110 (2016)]. Moreover, the enhancement was not observed in a quasi-one-dimensional limit of the Coulomb ground state at $\nu{=}1/2$ [Geraedts et al., Science 352, 197 (2016)]. Here, we revisit the problem of entanglement scaling in the CFL state realized in a two-dimensional continuum system. Using Monte Carlo evaluation of the second Rényi entropy $S_{2}$ for the CFL variational wave function, we show that the entanglement enhancement is present not only at $\nu{=}1/2$ but also at $\nu{=}1/4$, as well as in bosonic CFL states at $\nu{=}1$ and $\nu{=}1/3$ fillings. In all cases, we find the scaling of $S_{2}$ with subsystem size to be enhanced compared to the non-interacting case, and insensitive to the choice of geometry and projection to the lowest Landau level. We also demonstrate that the variance of the particle number in a subsystem obeys area-law scaling with a universal subleading corner contribution, in stark contrast with free fermions. Our results establish the enhanced entanglement scaling and suppressed charge fluctuations as fingerprints of non-Fermi-liquid correlations in the CFL state.
@@
Introduction.—The fundamental quasiparticles of fractional quantum Hall states are composite fermions (CFs) – electrons dressed by an even number of vortices of the magnetic flux [1]. At even-denominator fillings of the lowest Landau level (LLL), CFs can form a compressible state – the “composite fermion Fermi liquid” (CFL) [2, 3, 4], which has been observed in numerous experiments [5]. The CFL state has also been argued to emerge in systems with flat bands in the absence of a magnetic field, such as twisted moiré materials [6, 7].
@@
Flux attachment – the mechanism behind the formation of CFs [8, 2, 9] – couples their Fermi surface with an internal statistical U(1) gauge field, placing CFLs inside a larger class of “non-Fermi liquids” (NFLs) (see Ref. [10] for a recent review). On the one hand, the CFL is a pristine example of a stable NFL phase that emerges solely due to electron-electron interactions. On the other hand, the existence of well-defined quasiparticles implies there may be a good degree of qualitative similarity between CFL and conventional Fermi liquids [11, 12, 13]. Indeed, the Fermi wave vector of CFs is consistent with that derived from the electron density [14, 15, 16] and it also satisfies the Luttinger theorem [17, 18]. Thus, pinpointing the NFL nature of the CFL is a nontrivial task.
@@
One sensitive diagnostic of any quantum state is its bipartite entanglement entropy. The entanglement entropy scaling in non-interacting Fermi liquids obeys the well-known Widom formula [19, 20, 21], wherein the “area law” – ubiquitous to gapped systems – gets modified by a multiplicative logarithmic correction. It has been argued that turning on interactions leads to the same scaling [22], leaving open the value of the prefactor [23, 24]. Previous numerical tests of this hypothesis for the CFL state have arrived at conflicting conclusions. Infinite density matrix renormalization group (iDMRG) study of the Coulomb ground state at $\nu{=}1/2$ revealed no sign of correction to the Widom formula [25]. For the continuum CFL variational wave function of Rezayi and Read [26], Ref. [27] found a significant multiplicative prefactor to the entanglement entropy scaling (${\approx}2$ in system sizes on the order of $40$ electrons). Lastly, Ref. [28] also found an enhancement for a lattice analog of the CFL wave function. However, based on the scaling of different contributions to the entanglement from the sign and modulus of the wave function, Ref. [28] ruled out the enhancement as a finite-size effect. These studies therefore raise an important question: is there any distinctive entanglement signature of the CFL state in the thermodynamic limit?
Figure 1: (a): Illustration of the Monte Carlo SWAP algorithm: interchanging the highlighted regions of two copies of the system yields the second Rényi entropy $S_{2}$. For entanglement calculations, we use circular subregions, while for charge fluctuations we use the square (dashed) subregion to capture the corner contributions. (b)-(c): $S_{2}$ of CFLs on the square torus for projected (b) and unprojected (c) wave functions at fillings $\nu{=}1/m$, with $m{=}1,2,3,4$. For all fillings, $S_{2}$ exhibits a multiplicative correction ${\sim}1.7$ to the Widom formula [Eq. (3)]. The $2k_{F}$ Friedel oscillations [2] are visible in all cases. All data is for $N{=}37$ particles. Crosses represent Monte Carlo data, while the continuous line is the exact diagonalization result for free fermions. (d) Extrapolations of the ratio between the slope $\kappa$ ($S_{2}$ vs. $\lambda\ln\lambda$) of the CFL at fillings $\nu{=}1/2$ and $\nu{=}1$ and the free-fermion slope $\kappa_{\mathrm{w}}$. We used system sizes $N{=}12{-}60$ and data points in the range which yield the extrapolated free-fermion slope ${\approx}0.25$ [29].
@@
Here we report a systematic study of entanglement in several spinless CFLs described by the Rezayi-Read continuum wave function [26]. We compare the fermionic CFLs at filling factors $\nu{=}1/2$ and $\nu{=}1/4$, which are distinguished by the number of vortices attached to each electron (two and four, respectively), and contrast those against an odd number of vortices in bosonic CFLs at $\nu{=}1$ and $\nu{=}1/3$ fillings. We employ Monte Carlo techniques to evaluate the second Rényi entropy using the SWAP algorithm [30]. In all CFL states considered, we find an enhancement of entanglement scaling, consistent with Ref. [27]. At the same time, we reconcile these findings with entropy decompositions considered in Ref. [28] by showing that the latter is more sensitive to finite-size effects and LLL projection compared to the total entropy. We demonstrate the robustness of our conclusions in sphere and torus geometries. Finally, we use the same setup to evaluate charge fluctuations in a subsystem, for which analytical predictions are available [31]. We confirm that charge fluctuations in the CFL state follow the expected area-law scaling, in sharp contrast with free fermions. Moreover, within the same range of system sizes, we accurately extract the subleading corner contribution to the charge fluctuations, in agreement with the expected “super-universal” value [31] that also holds for gapped [32, 33] and critical states [34, 35] in (2+1)-dimension.
@@
CFL wave function.—The CFL state at filling factor $\nu{=}1/m$ is described by the ansatz wave function [26]:
$\Psi_{m}^{\mathrm{CFL}}=\mathrm{Det}\big{[}\chi_{n}(z_{j},z^{*}_{j})\big{]}{\prod}_{i{<}j}(z_{i}{-}z_{j})^{m}e^{-\frac{1}{4}\sum_{k}|z_{k}|^{2}}.$ (1)
@@
The first term is a Slater determinant of single-particle orbitals $\chi_{n}$ occupied by CFs in zero magnetic field, and the remainder is the Laughlin wave function at filling $\nu{=}1/m$ [36], expressed in terms of particle coordinates $z_{j}{=}x_{j}{+}iy_{j}$ with the magnetic length set to unity. Note that, due to the antisymmetry of the Slater determinant, these wave functions describe fermionic CFL states for $m$ even and bosonic CFL states for $m$ odd. The wave functions of Eq (1) can be adapted to the spherical and torus geometries considered in this work, see Refs. [26, 37, 38] and the Supplementary Material (SM) [29]. Moreover, the wave function of Eq. (1) is not constrained to reside in the LLL. To make contact with previous work, we will project the wave function to the LLL using the Jain-Kamilla method [39, 27, 40, 41]. Nevertheless, we will probe the sensitivity of our results by also considering the unprojected wave function as written in Eq. (1).
@@
Rényi entropy.— The object of our study is the second Rényi entropy of the reduced density matrix $\hat{\rho}_{A}$ of region $A$:
@@
obtained by tracing over the complement $\bar{A}$. In most of our calculations, we take $A$ to be a circular subregion of radius $r_{A}$, see Fig. 1(a), although later we will also consider a square-shaped region to determine corner contributions to charge fluctuations in $A$. To extract the scaling, we fix a large system size and vary $r_{A}$. $S_{2}$ can be conveniently evaluated via variational Monte Carlo by replacing the trace with an expectation value of a SWAP operator between two copies of the system [30], which we review in the SM [29].
For free fermions in 2D, the area-law scaling of $S_{2}$ is violated by a term that depends logarithmically on the dimensionless quantity $\lambda{\equiv}k_{F}r_{A}$, with $k_{F}$ denoting the Fermi momentum. The violation is given by the asymptotic Widom formula [19, 20, 21]:
@@
where $\kappa_{\mathrm{w}}$ is the Widom coefficient that depends solely on the geometry and effective central charge of a (1+1)-dimensional chiral relativistic fermion [42, 43]. For a circular 2D Fermi surface, we have $\kappa_{\mathrm{w}}{=}1/4$. One of the goals of this paper is to ascertain whether the entropy of CFL states follows the scaling $S_{2}{=}\kappa\lambda\ln\lambda+\ldots$, analogous to Eq. (3), and whether $\kappa$ is the same as $\kappa_{\mathrm{w}}$.

### v2

# Entanglement scaling and charge fluctuations in a Fermi liquid of composite fermions

@@ Abstract
The composite fermion Fermi liquid (CFL) state at $\nu=1/2$ filling of a Landau level is a paradigmatic non-Fermi liquid borne out purely by Coulomb interactions. But in what ways is this exotic state of matter different from a Fermi liquid? The CFL entanglement entropy was indeed found to exhibit a significant enhancement compared to free electrons [Shao et al., Phys. Rev. Lett. 114, 206402 (2015)], which was subsequently ruled out as a finite-size effect by the study of a lattice CFL analog [Mishmash and Motrunich, Phys. Rev. B 94, 081110 (2016)]. Moreover, the enhancement was not observed in a quasi-one-dimensional limit of the Coulomb ground state at $\nu=1/2$ [Geraedts et al., Science 352, 197 (2016)]. Here, we revisit the problem of entanglement scaling in the CFL state realized in a two-dimensional electron gas. Using Monte Carlo evaluation of the second Rényi entropy $S_{2}$ for the CFL variational wave function, we show that the entanglement enhancement is present not only at $\nu=1/2$ but also at $\nu=1/4$, as well as in bosonic CFL states at $\nu=1$ and $\nu=1/3$ fillings. In all cases, we find the scaling of $S_{2}$ with subsystem size to be enhanced compared to the non-interacting case, and insensitive to the choice of geometry and projection to the lowest Landau level. We also demonstrate that, for CFL states, the variance of the particle number in a subsystem obeys area-law scaling with a universal subleading corner contribution, in stark contrast with free fermions. Our results establish the enhanced entanglement scaling and suppressed charge fluctuations as fingerprints of non-Fermi-liquid correlations in CFL states.
I Introduction
@@
The fundamental quasiparticles of fractional quantum Hall states are composite fermions (CFs) – electrons dressed by an even number of vortices [1]. At even-denominator fillings of the lowest Landau level (LLL), the CFs can form a compressible state – the “composite fermion Fermi liquid” (CFL) [2, 3, 4], which has been observed in numerous experiments [5]. The CFL state has also been argued to emerge in systems with flat bands in the absence of a magnetic field, such as twisted moiré materials [6, 7].
@@
Flux attachment – the mechanism behind the formation of CFs [8, 2, 9] – couples the Fermi surface of CFs with an internal statistical U(1) gauge field, placing the CFLs inside a larger class of “non-Fermi liquids” (NFLs) (see Ref. [10] for a recent review). On the one hand, the CFL is a pristine example of a stable NFL phase that emerges solely due to electron-electron interactions. On the other hand, the existence of well-defined CF quasiparticles implies there may be a qualitative similarity between CFL and conventional Fermi liquids [11, 12, 13]. Indeed, the Fermi wave vector of CFs is consistent with that derived from the electron density [14, 15, 16] and it also satisfies the Luttinger theorem [17, 18]. Thus, pinpointing the NFL nature of the CFL is a nontrivial task.
@@
One sensitive diagnostic of a quantum state is its bipartite entanglement entropy. The entanglement entropy scaling in non-interacting Fermi liquids obeys the well-known Widom formula [19, 20, 21], wherein the “area law” – ubiquitous to gapped systems – gets modified by a multiplicative logarithmic correction. It has been argued that turning on interactions does not modify the scaling [22], leaving open the value of the prefactor [23, 24]. Previous numerical tests of this hypothesis for the CFL state have arrived at conflicting conclusions. Infinite density matrix renormalization group (iDMRG) study of the Coulomb ground state at $\nu=1/2$ revealed no sign of correction to the Widom formula [25]. For the continuum CFL variational wave function of Rezayi and Read [26], Ref. [27] found a significant multiplicative prefactor to the entanglement entropy scaling ($\approx 2$ in system sizes on the order of $40$ electrons). Lastly, Ref. [28] also found an enhancement for a lattice analog of the CFL wave function. However, based on the scaling of different contributions to the entanglement from the sign and modulus of the wave function, Ref. [28] ruled out the enhancement as a finite-size effect. These studies therefore raise an important question: is there any distinctive entanglement signature of the CFL state in the thermodynamic limit?
@@
Here we report a systematic study of entanglement in several CFLs described by the Rezayi-Read wave function [26], known to have a large overlap with the Coulomb ground state [29, 30]. We assume full spin polarization and compare the fermionic CFLs at fillings $\nu=1/2$ and $\nu=1/4$, which are distinguished by the number of vortices attached to each electron (two and four, respectively), and contrast those against an odd number of vortices in bosonic CFLs at $\nu=1$ and $\nu=1/3$ fillings. We employ Monte Carlo techniques to evaluate the second Rényi entropy using the SWAP algorithm [31]. In all CFL states considered, we find a pronounced enhancement of entanglement scaling, consistent with Ref. [27], up to the largest system sizes accessible ($N\lesssim 60$). At the same time, we reconcile these findings with entropy decompositions considered in Ref. [28] by showing that the latter is more sensitive to finite-size effects and LLL projection compared to the total entropy. We demonstrate the robustness of our conclusions in sphere and torus geometries. Finally, we use the same setup to evaluate charge fluctuations in a subsystem, for which analytical predictions are available [32]. We confirm that charge fluctuations in the CFL state follow the expected area-law scaling, in sharp contrast with free fermions. Moreover, within the same range of system sizes, we accurately extract the subleading corner contribution to the charge fluctuations, in agreement with the expected “super-universal” value [32] that also holds for gapped [33, 34] and critical states [35, 36] in (2+1)-dimension.
The remainder of this paper is organized as follows. In Sec. II we introduce the CFL wave functions on the torus and sphere geometries and evaluate their entanglement entropy. Section III contains the results for charge fluctuations in the same CFL states. In Sec. IV we discuss the scaling of entanglement entropy in the context of modulus and sign entropy decompositions, carefully contrasting the results against free fermions. Our conclusions and a discussion of open questions are presented in Sec. V, while Appendices contain further technical details, including a description of the Monte Carlo algorithm and the LLL projection of the wave functions.
II The CFL wave function and its Rényi entropy
@@
The essential physics of the CFL state at filling factor $\nu=1/m$ is captured by the ansatz wave function [26, 2]:
$\Psi_{m}^{\mathrm{CFL}}(z_{1},\ldots,z_{N})=\mathrm{Det}\big{[}\chi_{n}(z_{j},z^{*}_{j})\big{]}\psi_{m}^{\mathrm{L}}(z_{1},\ldots,z_{N}),$ (1)
@@
expressed in terms of particle coordinates $z_{j}=x_{j}+iy_{j}$. The first term in Eq. (1) is a Slater determinant of single-particle orbitals $\chi_{n}$ occupied by CFs in zero magnetic field, while the second is the Laughlin-Jastrow wave function at filling factor $\nu=1/m$ [37]. Due to the antisymmetry of the Slater determinant, these wave functions describe fermionic CFL states for $m$ even and bosonic CFL states for $m$ odd. As discussed below, the wave functions of Eq. (1) can be adapted to the spherical and toroidal geometries considered in this work, see Refs. [26, 38, 39]. Moreover, the wave function of Eq. (1) is not constrained to reside in the LLL. Therefore, to make contact with previous work, we will project the wave function to the LLL using the Jain-Kamilla (JK) method [40, 27, 41, 42]. Nevertheless, we will probe the sensitivity of our results by also studying the unprojected wave function as in Eq. (1).
Figure 1: (a)-(b): $S_{2}$ Rényi entropy for projected (a) and unprojected (b) CFL wave functions at fillings $\nu=1/m$, with $m=1,2,3,4$. For all fillings, $S_{2}$ exhibits a multiplicative correction $\sim 1.7$ to the Widom formula, Eq. (4). The $2k_{F}$ Friedel oscillations [2] are visible in all cases. All data is for $N{=}37$ particles on a square torus. The symbols represent Monte Carlo data, while the continuous black line is the exact diagonalization result for free fermions. (c) Extrapolations of the ratio between the slope $\kappa$ ($S_{2}$ vs. $\lambda\ln\lambda$) of the CFL at fillings $\nu=1/2$ and $\nu=1$ and the free-fermion slope $\kappa_{\mathrm{w}}$. We used system sizes $N=12-60$ and data points in the range which yield the extrapolated free-fermion slope $\approx 0.25$, see Appendix D for details.
@@
The central object of our study is the second Rényi entropy of the reduced density matrix $\hat{\rho}_{A}$ of region $A$:
@@
obtained by tracing over the complement $\bar{A}$ of region $A$. We typically take $A$ to be a circle of radius $r_{A}$, although in Sec. III we will also consider a square-shaped region to determine corner contributions to charge fluctuations in $A$. To extract the scaling, we fix a large system size and vary $r_{A}$. $S_{2}$ can be conveniently evaluated via variational Monte Carlo by replacing the trace with an expectation value of a SWAP operator between two copies of the system [31], which is reviewed in Appendix A.
For any wave function $\Psi$ that can be written as a Slater determinant, $\Psi={\rm Det}[\phi_{m}(\mbox{\boldmath$r$}_{n})]$, where the single-particle orbitals $\phi_{m}$ are orthonormalized over the full space, the entanglement entropy can be efficiently computed using the correlation matrix $\mathcal{A}$ [43]. The matrix elements of $\mathcal{A}$ are given by the overlap integrals on the subsystem, $\mathcal{A}_{mn}=\int_{A}\mathrm{d}\mathbf{r}\,\phi_{m}^{*}(\mathbf{r})\phi_{n}(\mathbf{r})$. Denoting the eigenvalues of $\mathcal{A}$ by $a_{m}$, the second Rényi entropy is given by [44, 45]
$S_{2}^{\mathrm{free}}=-\sum_{m}\ln\left(a_{m}^{2}+(1-a_{m})^{2}\right).$ (3)
Explicit expressions for $\mathcal{A}_{mn}$ in different geometries considered in this work are given in Appendix B. With this, one can use Eq. (3) to efficiently evaluate the Rényi entropy of free fermions in systems of finite size.
In the asymptotic limit, for free fermions in 2D, the area-law scaling of $S_{2}$ is violated by a term that depends logarithmically on the dimensionless quantity $\lambda\equiv k_{F}r_{A}$, with $k_{F}$ denoting the Fermi momentum. The violation is given by the Widom formula [19, 20, 21]:
@@
where $\kappa_{\mathrm{w}}$ is the Widom coefficient that depends solely on the geometry and effective central charge of a (1+1)-dimensional chiral relativistic fermion [46, 47]. For a circular 2D Fermi surface, we have $\kappa_{\mathrm{w}}=1/4$. One of the goals of this paper is to ascertain whether the entropy of CFL states follows the scaling $S_{2}=\kappa\lambda\ln\lambda+\ldots$, analogous to Eq. (4), and whether $\kappa$ is the same as $\kappa_{\mathrm{w}}$. In the remainder of this section, we directly evaluate the $S_{2}$ Rényi entropy for the wave functions in Eq. (1) on the torus and sphere, and compare them against Eq. (4).
II.1 Torus
On the torus, the CFL state at filling factor $\nu=1/m$ is described by the wave function in Eq. (1), where we take the single-particle orbitals to be plane waves, $\chi_{n}(\mathbf{r}_{j})=\exp(i\mathbf{k}_{n}\cdot\mathbf{r}_{j})$, with $\mathbf{k}_{n}$ the wave vectors of the occupied states of the Fermi sea [26]. The latter can be determined from an empirical rule given in Refs. [39, 38, 27]. On the other hand, the Laughlin-Jastrow part of the wave function $\psi^{\mathrm{L}}_{m}$ has a more complicated form given by
$\psi^{\mathrm{L}}_{m}$ $=$ $\vartheta\begin{bmatrix}\frac{N-1}{2}\\ \frac{m(N-1)}{2}\end{bmatrix}\bigg{(}\frac{m\sum z_{i}}{L_{1}}\bigg{|}m\tau\bigg{)}$
$\prod_{i<j}\bigg{\{}\vartheta\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}\bigg{(}\frac{z_{i}-z_{j}}{L_{1}}\bigg{|}\tau\bigg{)}\bigg{\}}^{m}\exp\bigg{(}{\sum_{i}\frac{z_{i}^{2}-|z_{i}^{2}|}{4\ell^{2}}}\bigg{)},$
where the torus periodicity is ensured by the Jacobi $\vartheta$ functions with rational characteristics [48, 49, 50]. The Jacobi $\vartheta$ functions explicitly depend on the modular parameter $\tau$ of the torus and one of its dimensions, $L_{1}$, where $\ell$ denotes the magnetic length. The $\vartheta$ function in the first line of Eq. (II.1) describes the center of mass of the particles, and we have implicitly chosen it to carry zero momentum and obey strictly periodic conditions for the torus [51]. The remaining terms in Eq. (II.1) represent a periodized version of the familiar Laughlin-Jastrow factor in the infinite disk geometry [37].
While the wave function $\psi_{m}^{\mathrm{L}}$ in Eq. (II.1) is fully within the LLL, the multiplication via plane waves $\chi_{n}(\mathbf{r}_{j})$ brings the total wave function $\Psi^{\mathrm{CFL}}_{m}$ outside the LLL. Following Ref. [52], we obtain the family of LLL-projected CFL wave functions:
$\Psi^{\mathrm{CFL,proj[\bm{\alpha}]}}_{m}$ $=\vartheta\begin{bmatrix}\frac{N-1}{2}\\ \frac{m(N-1)}{2}\end{bmatrix}\bigg{(}\frac{m\left(\sum z_{i}+i\ell^{2}\sum k_{i}\right)}{L_{1}}\bigg{|}m\tau\bigg{)}$
×$\times\mathrm{Det}[g_{nl}^{[\bm{\alpha}]}]\exp\bigg{(}{\sum_{i}\frac{z_{i}^{2}-|z_{i}^{2}|}{4\ell^{2}}}\bigg{)}\,,$ (6)
where the key consequence of LLL projection are the CF “single-particle” orbitals $g_{nl}^{\bm{\alpha}}$:
$g_{nl}^{[\bm{\alpha}]}$ $=e^{k_{n}(k_{n}+2\overline{k}_{n})\ell^{2}/4}e^{i(k_{n}+\overline{k}_{n})z_{l}/2}$
×$\times\prod_{p=1}^{m/2}\prod_{j\neq l}\vartheta\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}\bigg{(}\frac{z_{l}+i\alpha_{p}k_{n}\ell^{2}-z_{j}}{L_{1}}\bigg{|}\tau\bigg{)},$ (7)
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{m/2})$ are the corresponding JK projection coefficients. Torus periodicity enforces the constraint $\sum_{p=1}^{m/2}\alpha_{p}=m$.
Note that $g_{nl}^{\bm{\alpha}}$ in Eq. (II.1) depend on the coordinates of all the particles, hence they are complicated many-body functions that only formally resemble single-particle orbitals. Essentially, the LLL projection shifts the attached vortices by an amount proportional to the wave number of the CF. The scheme in Eq. (II.1) encompasses fermionic CFL wave functions with even $m$; for bosons, one needs to attach one additional set of Jastrows outside the wave function for $m\geq 3$ and remove one for $m=1$. In Appendix C we verify the consistency of our results for different choices of $\bm{\alpha}$, showing that, in general, our results are insensitive to the details of the projection.
In Fig. 1(a)-(b) we show the evaluated $S_{2}$ for bosonic and fermionic CFL states at fillings $\nu=1/m$, with $m{=}1,2,3,4$. Based on our free-fermion benchmarks presented in Appendix D, to minimize the finite-size effect, we harvest data points such that the subsystem area ranges from 1% to 30% of the total area and use a fixed Fermi momentum of $k_{F}=\sqrt{2\nu}$. The extrapolated slope $\kappa$, in units of the Widom slope $\kappa_{\mathrm{W}}$, is plotted in Fig. 1(c) as a function of inverse system size.